OpenStudy (anonymous):

What do I need to do to the equation sqrt(x^2-x+1)=2sin(pix) to be able to apply Newton's Method?

OpenStudy (phi):

write it as \[f(x) = 2 \sin(\pi x) - (x^2-x+1)^{\frac{1}{2}} \] and solve using \[ x_{n+1} = x_n - \frac{f(x)}{f'(x)} \] you will need a rough idea of where the root(s) are. i.e a guess for x0

OpenStudy (phi):

x^2-x+1 is a "cup shaped" parabola with a minimum value at -b/(2a) = 1/2 and so sqrt( ) will have a min value at 0.707 so if it intersects the sin, it will be when the sin is positive, between 0 and pi I would use x0=0 and x0= pi as the starting values, and expect to find 2 roots.

OpenStudy (anonymous):

Okay. Thank you very much for the help!